OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x + sin x

OpenStudy (anonymous):

\(\cos 2x = \cos^2x-\sin^2x\), so \(\cos 2x+\sin x = \cos^2 x-\sin^2x+\sin x\).

OpenStudy (anonymous):

1 + 3 sin^2x

OpenStudy (anonymous):

What do you want done with that?

OpenStudy (anonymous):

no thats my anserw :) is that right

OpenStudy (anonymous):

No. How are you getting that? See my original response.

OpenStudy (anonymous):

1 - 2 sin2x + sin x

OpenStudy (anonymous):

Yep, that works (or \(2\,\cos^2 x - 1+\sin x\)).

OpenStudy (anonymous):

1 + 2 sin^2x + sin x could it be this one too?

OpenStudy (anonymous):

No, you need the minus sign there. \(1-2\,\sin^2 x +\sin x\) is totally fine.

OpenStudy (anonymous):

thanks :))